### v3

# Pushing the Limits of Sparsity: A Bag of Tricks for Extreme Pruning

Abstract
Pruning of deep neural networks has been an effective technique for reducing model size while preserving most of the performance of dense networks, crucial for deploying models on memory and power-constrained devices. While recent sparse learning methods have shown promising performance up to moderate sparsity levels such as 95% and 98%, accuracy quickly deteriorates when pushing sparsities to extreme levels due to unique challenges such as fragile gradient flow. In this work, we explore network performance beyond the commonly studied sparsities, and propose a collection of techniques that enable the continuous learning of networks without accuracy collapse even at extreme sparsities, including 99.90%, 99.95% and 99.99% on ResNet architectures. Our approach combines 1) Dynamic ReLU phasing, where DyReLU initially allows for richer parameter exploration before being gradually replaced by standard ReLU, 2) weight sharing which reuses parameters within a residual layer while maintaining the same number of learnable parameters, and 3) cyclic sparsity, where both sparsity levels and sparsity patterns evolve dynamically throughout training to better encourage parameter exploration. We evaluate our method, which we term Extreme Adaptive Sparse Training (EAST) at extreme sparsities using ResNet-34 and ResNet-50 on CIFAR-10, CIFAR-100, and ImageNet, achieving significant performance improvements over state-of-the-art methods we compared with.
1 Introduction
Network pruning [9, 10, 16, 22, 19, 15] is a widely-used technique for reducing a network’s parameters and compressing its size. Reducing model sizes is crucial for deploying models on edge devices with limited resources. Conventionally, pruning methods have focused on reducing parameters from pre-trained models. However, it requires at least as much computation as training a dense model as it must converge before pruning takes place. The Lottery Ticket Hypothesis [6, 7, 23] gives theoretical foundation that subnetworks have the potential to reach full performance even when trained from an initially sparse state. This insight has recently gained much traction in sparse training, a paradigm in which sparse networks are trained from scratch without the need for dense pre-training.
Sparse training methods can be broadly classified into two categories: static sparse training (SST), also sometimes referred to as Pruning at Initialization (PaI), where the sparsity pattern is pre-determined at initialization and remains fixed throughout training [17, 33, 30, 2], and dynamic sparse training (DST), where the sparsity pattern continuously evolves during training [25, 3, 4, 34]. These methods perform well at moderate sparsity commonly up to 98%, but experience rapid accuracy degradation beyond this due to limitations such as layer collapse and gradient issues [31, 33]. There is an increasing demand to push towards ’extreme sparsities’, particularly in resource-constrained settings where substantial computational and memory benefits may justify some accuracy trade-offs. Our work is motivated by this practical consideration, and thus we emphasize that our primary focus is on the deployable model used for inference rather than training efficiency. Despite this need, extreme sparsities remain relatively understudied, with only a few methods attempting to maintain usable accuracy at these extreme levels [31, 2, 28].
Figure 1: Illustration of EAST (top). Starting with ERK-initialized sparse network $\mathcal{A}$ = $\theta_{0}$, $\mathcal{M}_{0}$} at sparsity s, EAST employs three key components: DyReLU phasing, weight sharing, and cyclic sparsity to transform the network to final state $\mathcal{A}^{\prime}$ = $\theta_{T}$, $\mathcal{M}_{T}$}, achieving meaningful performance at extreme sparsity levels. The topology update box (bottom) illustrates the connectivity change throughout training: connections are first grown, then pruned, and eventually maintained with a fixed sparsity update schedule until completion.
Our work proposes a collection of adaptive methods that when combined can achieve meaningful performance at extreme sparsities. By leveraging three core strategies - 1) phased Dynamic ReLU activation, 2) weight sharing, 3) cyclic sparsity scheduling - we present a flexible framework that is capable of optimizing performance at extreme sparsities. Each component independently enhances the sparse model’s capacity as seen in our ablation studies, where individual strategy improves performance on their own. However, when used together, these methods can cohesively achieve an even stronger performance. This framework, which we term Extreme Adaptive Sparse Training (EAST), offers a modular approach that can be adapted to existing DST frameworks, effectively postponing their total performance collapse, as shown in Figure 2. Through empirical analysis, we find that EAST prevents gradient vanishing and encourages parameter exploration, achieving substantial and meaningful performance gain over our baselines. Our contribution and key achievements can be summarized as follows:
  • We introduce EAST - a method that combines phased Dynamic ReLU activation, weight sharing, and cyclic sparsity scheduling, aimed at retaining performance past common sparsity thresholds.
  • We perform extensive empirical evaluation of EAST on image classification tasks, and demonstrate that it can achieve superior performance at sparsities beyond 99.90%, surpassing all previous methods tested.
  • We explore and demonstrate that DyReLU when used as a drop-in replacement for ReLU at initialization at then replaced completely can still improve the parameter exploration at extreme sparsities, leading to higher performance than sparse models trained using ReLU only.
  • To the best of our knowledge, EAST is the first DST method to dynamically adjust both sparsity patterns and sparsity levels throughout training (as opposed to just the patterns). Our cyclic sparsity schedule allows models to explore parameters and stabilize at extreme sparsity, as evidenced by our ablation studies.
  • Unlike DCTpS, EAST achieves extreme sparsity without requiring dense matrix computations for inference, resulting in a truly lightweight and efficient model for edge deployment. We also demonstrate the versatility of our method by combining with DCT layers on top of DST, consistently improving accuracy at the tested settings.
Figure 2: Comparison of test accuracies across different sparsities. Each point represents median accuracy over three runs with different seeds, and the shaded regions highlight the variability across runs.
2 Related Work
Dynamic Sparse Training. Early work by Mocanu et al. demonstrated the viability of DST through SET [25], which used magnitude pruning and random weight regrowth. Subsequent criteria for regrowing connections have been developed, such as using the momentum of parameters in SNFS [3] or their gradient in RigL [4] to determine the salience of deactivated connections. GraNet [20] adapted RigL’s approach by starting with half the parameters and gradually increasing to target sparsity instead of starting from the target sparsity. ITOP [21] provides a deeper understanding of pre-existing DST methods and enhances parameter exploration during sparse training. Supticket [34] further enhances the performance of DST by leveraging weight averaging at the late training stage. NeurRev [18] addressed the issue of dormant neurons by removing harmful negative weights. Recent advances include Top-KAST [13], which maintains an auxiliary set of weights for enhanced exploration, and BiDST [14] which formulates DST as a bi-level optimization problem for joint weight-mask optimization. MEST [36] introduces a memory-efficient sparse training framework for edge devices, while Chase [35] adapts dynamic unstructured sparsity into hardware-friendly channel-wise sparsity, both optimizing performance on resource-constrained systems without sacrificing accuracy.
Extreme Sparsity Methods. Most current methods focus on sparsity levels at which pruned networks match or closely approximate the performance of their dense counterparts, a concept referred to as matching sparsity [8]. However, there is growing interest in extreme sparsities, where accuracy trade-offs become inevitable but potentially worthwhile for substantial computational gains. Yet there remains challenges, such as maintaining stable gradient flow and preventing layer collapse, which conventional pruning methods struggle to address. GraSP [33] and SynFlow [31] specifically addressed the gradient preservation and layer collapse issues at high sparsities via careful network initialization. SNIP [17] and SNIP-it [32] proposed PaI methods that showed promising results at higher sparsities, though primarily focused on static masks. De Jorge et al. [2] demonstrated that methods like SNIP and GraSP can perform worse than random pruning at 99% sparsity and beyond, and proposed iterative pruning approaches (Iter-SNIP and FORCE) that maintained meaningful performance up to 99.5% sparsity through gradual parameter elimination. DCTpS [28] introduced ”DCT plus Sparse” layers that combined a fixed dense DCT offset matrix with trainable sparse parameters to maintain propagation at extreme sparsities up to 99.99%.
Activation Functions. Activation functions in sparse networks have traditionally relied on static activation functions such as ReLU [27], which applies a fixed threshold to the input. Parametric variants such as PReLU [11] introduced learnable parameters to adapt the negative slope, offering slight improvements in gradient flow. Swish [29] and Mish [24] aim to smooth gradients for better optimization but remain static throughout training. Dynamic ReLU (DyReLU) [1] introduced an adaptive activation that dynamically adjusts the slopes of a ReLU based on the input.
3 Methodology
3.1 Activation Function
Figure 3: Positive pre-activations analysis in ResNet-34 at 99.99% sparsity. The left figure shows layerwise-comparison of positive pre-activations after DyReLU is completely converted to ReLU. The right figure shows their overall amount before, during and after DyReLU phasing.
In this work, when initializing a network, we replace the standard ReLU activation with a non-ReLU activation. We tested several activations and found that DyReLU worked the best, specifically the DyReLU-B variant, which dynamically adapts and adjusts channel-wise activation coefficients during training, allowing for richer expressivity during early training. As training progresses, we gradually transition from DyReLU to standard ReLU, eliminating the additional parameters. Preliminary details about DST, DyReLU and details on phasing out DyReLU can be found in the Supplementary Materials.
Understanding Gradient Flow Through Neuron Activity. With ReLU, non-positive pre-activations would result in zero gradient and prevent the learning signal with ReLU. To better understand why DyReLU phasing improves performance at extreme sparsity, we analyzed pre-activation values throughout training. As shown in Figure 3, the model trained solely with ReLU maintains around 1.3% positive pre-activations throughout training, with many deeper layers showing none, interrupting gradient propagation.
In contrast, our DyReLU phasing approach naturally begins from a high 56.8%, decreases during the transition phase, and stabilizes at 4.5% even after DyReLU is completely converted to ReLU, three times higher than with ReLU alone. The deeper layers also retain some positive pre-activations, preserving critical activation pathways that would otherwise be lost. This allows extremely sparse networks to continue learning where conventional approaches would struggle due to interrupted gradient propagation. We detail the ablation studies on the effect of DyReLU phasing in section 4.2.
3.2 Weight Sharing
Much research in sparse training has been about mask optimization - determining the best subset of parameters to keep. However, at extreme sparsities, gradient flow can collapase even with an optimal parameter subset. Therefore, we ask a different question: Can we increase the number of parameter paths in the computational graph while maintaining the same number of learnable parameters?
Our weight sharing mechanism strategically reuses parameters within each layer’s residual blocks, allowing the remaining active parameters to participate in multiple paths of the computational graph. If at a certain sparsity, we have $n$ remaining parameters in the original model, we keep exactly this many unique parameters to learn, but some of these now appear in multiple locations in the computational graph, multiplying their contribution without increasing storage requirements. During backpropagation, shared parameters recieve gradient signals from multiple positions, strengthening gradient flow.
Figure 4: A layer in ResNet-34 with 4 blocks. Block 3 and Block 4 share parameters (and the masks) with block 2. Specifically, conv1 layer (green) of block 3 reuses conv1 layer of block 2, and is multiplied by a learnable scaling factor; similarly, its conv2 layer (blue) reuses conv2 layer of block 2, and is multiplied by another scaling factor.
Specifically, we choose a block inside a layer and share its parameters (along with their masks) with all blocks after it. These later blocks reference the single block tensor that they share with, and multiply by a learnable scaling factor for forward and backward passes. For example, in a ResNet layer with 4 blocks (Figure 4), if the second block shares parameters with blocks 3 and 4, a parameter in block 2 effectively appears three times in the computation graph while only being stored and optimized once.
Thus, we adjust sparsity calculation to account for weight sharing, such that sparsity $\text{sparsity s}=1-\frac{\lVert\mathcal{M}\odot\theta_{\text{learnable}}\rVert_{0}}{\lVert\theta\rVert_{0}}$, where $\lVert$.$\rVert$ . ∥ is the $\ell_{0}$-norm, $\theta_{\text{learnable}}$ is the unique learnable parameter count, and $\theta$ is the theoretical parameter count in the original network with a mask $\mathcal{M}\in\{0,1\}^{|\theta|}$. Formally, let $L$ be the number of residual blocks in a layer, and $R$ be the block that shares parameters ($1<R\leq L$), where the remaining $L-R$ blocks share parameters with the $R$-th block. Let $\theta_{i}$ represent the parameters of the $i$-th block. Our weight sharing scheme can be expressed as:
$\theta_{i}=\begin{cases}\theta_{i}&\text{if }i\leq R\\ \theta_{R}&\text{if }R<i\leq L\end{cases}$
3.3 Cyclic Sparsity
While standard DST methods such as SET and RigL adjust the sparsity pattern, sparsity $s$ is consistent throughout training - pruning and regrowing of weights take place at the same $t$. i.e., $\mathcal{M}(t)=\text{Prune}(\theta(t))\cup\text{Grow}(\nabla_{\theta}\mathcal{L})$. In our work, we introduce a cyclic sparsity schedule where the network’s sparsity level itself evolves dynamically during training. Instead of maintaining a constant sparsity, we cyclically adjust the sparsity $s(t)$ between a maximum value $s_{\text{max}}$ and a minimum value $s_{\text{min}}$. The sparsity level $s(t)$ is governed by a cosine schedule:
$s(t)=s_{\text{min}}+\frac{s_{\text{max}}-s_{\text{min}}}{2}\cdot\left(1+\cos\left(\frac{2\pi t}{T_{c}}\right)\right)$
where $T_{c}$ is the period of the cyclic sparsity. This schedule ensures that the sparsity grows smoothly to $s_{\text{max}}$ and then gradually reduces to $s_{\text{min}}$. As a result, the network experiences phases of higher and lower connectivity, promoting better parameter exploration during training. Once the cyclic phase ends at $T_{c}$, we fix the sparsity and switch to a static pruning and growth update. The details of our method are summarized in Algorithm 1.
Algorithm 1 Pseudocode for EAST
1:Sparse neural network with weights $\theta$, maximum sparsity $s_{\text{max}}$, minimum sparsity $s_{\text{min}}$, end of cyclic sparsity $T_{c}$, update frequency $\Delta T$, prune rate $r_{p}$, DyReLU phase start $T_{s}$, end $T_{e}$
2:Trained $\theta_{s}$
3:Initialize: $\theta_{s}$ at $s_{\text{max}}$, $\phi\leftarrow\text{DyReLU}$
4:for $t=1$ to $T_{\text{end}}$ do
5:     for $T_{s}\leq t\leq T_{e}$ do $\triangleright$ DyReLU to ReLU
6:         $\phi\leftarrow\beta\text{DyReLU}+(1-\beta$) ReLU
7:         Update($\beta$)
8:     end for
9:     if $t\bmod\Delta T==0$ then
10:         if $t<=T_{c}$ then
11:              $s_{\text{target}}\leftarrow s_{\text{min}}+\frac{s_{\text{max}}-s_{\text{min}}}{2}\cdot(1+\cos(2\pi t/T_{c}))$triangleright Update target sparsity
12:              if $s_{\text{target}}>s_{\text{current}}$ then
13:                  $\theta_{s}\leftarrow\theta_{s}-\text{ArgTopK}(-|\theta_{s}|,(s_{\text{target}}-$
14:                  $s_{\text{current}})\lVert\theta_{s}\rVert_{0})$triangleright Magnitude pruning
15:              else if $s_{\text{target}}<s_{\text{current}}$ then
16:                  $\theta_{s}\leftarrow\theta_{s}+\text{ArgTopK}(|\nabla_{\theta}\mathcal{L}|,(s_{\text{current}}$
17:                  $-s_{\text{target}})\lVert\theta_{s}\rVert_{0})$triangleright Gradient regrowth
18:              end if
19:         $\triangleright$ Switch to fixed update
20:              $\theta_{s}\leftarrow\theta_{s}-\text{ArgTopK}(-|\theta_{s}|,r_{p}\lVert\theta_{s}\rVert_{0})$
21:              $\theta_{s}\leftarrow\theta_{s}+\text{ArgTopK}(|\nabla_{\theta}\mathcal{L}|,r_{p}\lVert\theta_{s}\rVert_{0})$
22:         end if
23:     end if
24:end for
25:return $\theta_{s}$
4 Experiments
Evaluation Protocol. Our experiments include image classification using ResNet-34 and ResNet-50 [12] on the benchmark datasets CIFAR-10 and CIFAR-100, as well as ResNet-50 on ImageNet, following the same evaluation protocol as other SOTA methods in terms of datasets and backbones. We compare our method with DST methods SET[25] and RigL[4], and a competitive SST method SynFlow[31], which also experiments at very high sparsities. For SynFlow, we use the official repository with its default hyperparameters to evaluate its performance. For SET and RigL, we use the repository provided by ITOP [21]. We keep all hyperparameters as they are optimally configured in the paper, including $\Delta T$, ERK initialization, and the mask update interval of 1500 and 4000 for SET and RigL, respectively.
In addition, we benchmark DCT, the SOTA method specifically focused on extreme sparsity using the repository provided by the paper [28]. While this method performs very competitively at extreme sparsity settings, it comes at a cost due to the computation of a dense matrix matching the full architecture size, which adds an extra layer of computational overhead and requires extra memory. This method achieves the best performance when paired with RigL, which we evaluate, and then we combine it with our approach to evaluate if our method can improve it further. In this integration, we use only DyReLU phasing and cyclic density, leaving weight sharing to prevent conflicts with the DCT matrices embedded in the model architecture. Here, we use $\Delta T$ = 100 to match results from the original paper, and we also use this value for our integration to ensure fairness.
Method CIFAR-10 CIFAR-100
99% 99.90% 99.95% 99.99% 99% 99.90% 99.95% 99.99%
ResNet-34
Synflow 82.03 61.61 56.33 10.00 57.44 22.91 1.00 1.00
SET 93.09 82.70 70.84 10.00 71.97 47.64 32.14 1.00
RigL 92.92 85.71 81.47 10.03 71.72 50.26 40.92 1.14
EAST (ours) 93.03 86.76 83.83 70.57 72.01 52.16 42.53 18.84
DCTpS+RigL 89.04 83.96 80.80 77.38 63.12 51.30 45.68 37.59
DCTpS+RigL+EAST (ours) 89.22 84.19 81.96 77.54 64.33 53.21 47.80 38.02
ResNet-50
Synflow 77.06 48.62 10.00 10.00 48.91 12.07 9.40 1.00
SET 91.97 27.05 16.04 10.00 70.35 8.59 1.26 1.00
RigL 92.75 58.41 30.14 10.00 70.50 51.93 7.02 1.00
EAST (ours) 93.12 86.16 76.60 54.63 70.64 52.84 14.23 1.88
DCTpS+RigL 89.66 85.98 85.03 83.13 64.85 55.98 52.71 49.63
DCTpS+RigL+EAST (ours) 89.67 87.10 85.87 83.88 64.92 58.17 55.60 51.33
Table 1: Accuracy comparison on CIFAR-10 and CIFAR-100 at different sparsity levels. Top section compares sparse training methods at each sparsity level. Bottom section shows results with DCTpS, which achieves higher accuracies at $\geq$99.95sparsity, but requires dense matrix computations. The best results for each category are in bold.
Method Inference FLOP (M) Inference Network Size
99.95% 99.99% Comp. Cost Theoretical GPU-Supported GPU-Supported Params
ResNet-50 (Dense) 1297.83 - $\mathcal{O}(mn)$ $\mathcal{O}(N)$ $\mathcal{O}(N)$ 23.5M
SET 1.24 (0.001$\times$) 0.26 (0.0002$\times$) $\mathcal{O}(pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N)$ 23.5M
RigL 1.24 (0.001$\times$) 0.26 (0.0002$\times$) $\mathcal{O}(pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N)$ 23.5M
EAST (w/o WS) 1.24 (0.001$\times$) 0.26 (0.0002$\times$) $\mathcal{O}(pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N)$ 23.5M
EAST (w/ WS) 2.59 (0.002$\times$) 0.52 (0.0004$\times$) $\mathcal{O}(pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N-N_{s})$ 13.9M
DCTpS+RigL 40.26 (0.031$\times$) 39.09 (0.030$\times$) $\mathcal{O}(q\log q+pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N)$ 23.5M
DCTpS+RigL+EAST 40.26 (0.031$\times$) 39.09 (0.030$\times$) $\mathcal{O}(q\log q+pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N)$ 23.5M
Table 2: Comparison of computational complexity and network size of ResNet-50 on different methods. FLOPs are given alongside multiplicative change from dense model ($\times$). Let $N$ denote the total parameters in ResNet-50, $N_{s}$ the total shared parameters, $P\in(0,1)$ the global density, and $m\times n$ the size of flattened weight tensors with density $p$, where $q=\max(m,n)$. ”GPU-Supported Params” refers to the actual number of parameters on a commercial GPU without native support for irregular sparse patterns.
For all DST methods, we use SGD with momentum as our optimizer. The momentum coefficient is set to 0.9, and L2 regularization coefficient is set to 0.0001. We set the training epoch to 250 and use a learning rate scheduler that decreases from 0.1 by a factor of 10X at half way and three-quarters of way through training. For the DCT experiments, we train for 200 epochs with the Adam optimizer and a learning rate of 0.001. A more detailed description of the experimental setups and hyperparameters is included in the Supplementary Material. All experiments in our work are conducted on an A100 GPU.
CIFAR-10 and CIFAR-100. We evaluate EAST with ResNet-34 and ResNet-50 backbones, with a particular focus on extreme sparsities (99%-99.99%). We repeat our experiments three times and report the average accuracy in Table 1 (variance visualized in Figure 2), which presents results across different configurations.
At 99%, all dynamic training methods achieve comparable performance, significantly outperforming SynFlow. This is unsurprising as a static mask is expected to lead to worse performance than a dynamic mask [4, 26]. EAST’s advantages become increasingly pronounced at higher sparsities. At 99.90% and 99.95%, EAST beats the next best performing method RigL by a clear margin ranging between 1-2% with ResNet-34 on both datasets, showing significant gains in performance. The difference becomes even more clear with ResNet50, where other methods take a huge plummet on accuracy and EAST shows a more gradual decline. EAST displays an improvement of 38-40% over RigL on CIFAR-10, and 1-7% on CIFAR-100 at 99.90% and 99.95%. The most striking difference appears at 99.99%. We observe that all other methods result in complete collapse to random performance on either dataset and model. In contrast, EAST maintains meaningful accuracies except for ResNet50 on CIFAR-100. For instance, EAST achieves 70.57% on CIFAR-10 compared to 10% random baselines. This pattern holds across both models, though ResNet50 shows more degradation at the highest sparsity levels, presumably due to more prominent vanishing gradient problems in a deeper network, with even fewer parameters spread across each layer.
We evaluate DCTpS and find that while DCTpS demonstrates a clear advantage at 99.95% and 99.99%, it performs worse than SET, RigL and EAST at 99% and sometimes 99.9%. For instance, EAST achieves 3-8% higher accuracy than DCTpS at 99%, without involving additional overhead. To test the versatility of EAST, we further combine it with DCTpS and observe that EAST provides consistent improvements across all configurations tested.
99.50% Sparsity 99.90% Sparsity
Method Top-1 (%) Top-5 (%) Top-1 (%) Top-5 (%)
SynFlow 0.10 0.50 0.10 0.50
SET 0.10 0.50 0.10 0.50
RigL 32.50 58.58 8.53 24.18
DCT 0.10 0.50 0.10 0.50
DCT+RigL 19.99 43.34 6.62 18.84
EAST (ours) 34.45 62.44 11.23 29.47
Inference Performance
Throughput (imgs/sec) Latency (ms)
DCT+RigL 1229.25 10.18
EAST (ours) 1638.41 6.80
Table 3: Performance comparison on ImageNet at extreme sparsity levels using ResNet-50. For throughput and latency measurements, we use batch sizes of 2 and run 100 times and report the average.
ImageNet. Prior methods like SynFlow and DCTpS were evaluated primarily on smaller datasets like CIFAR and Tiny-ImageNet. We extend these baselines to ImageNet by implementing and adapting their published code and evaluate the scalability of EAST on ImageNet using ResNet-50 at sparsity 99.5% and 99.9%. Given the much shortened training epochs for ImageNet, we implement only the DyReLU phasing and weight sharing components, excluding cyclic sparsity. As presented in Table 3, EAST has demonstrated capability in maintaining network robustness on large-scale dataset, outperforming previous SOTA by a significant margin, achieving approximately 2% and 2.7% higher top-1 accuracy than RigL. This advantage is more notable when comparing to other methods including SynFlow and SET, which fail to maintain meaningful performance and collapase to random accuracy.
Interestingly, we observe different dynamics on ImageNet compared to CIFAR experiments. While DCTpS showed competitive advantages on smaller datasets, it does not exhibit the same competitive advantages as seen before, even with the extra overhead. DCTpS alone results in performance collapse and it performs worse than RigL. When combined with RigL, the accuracy improvements are modest and may be attributed primarily to RigL’s robust dynamic sparse exploration policy, rather than the DCT offset. On the inference efficiency front, EAST demonstrates practical advantages. Unlike DCTpS, which relies on a dense DCT matrix, EAST does not have such a drawback, as all DyReLU activations have been replaced with regular ReLU and thus runs just like the original ResNet model with ReLU. Therefore, EAST should theoretically have similar inference speed as the other DST methods, but faster than DCTpS for inference. To verify this, we ran both models for inference 100 times, taking their average and finding that DCTpS has 1.5$\times$ higher latency and 1.33$\times$ lower throughput than ours.
4.1 Complexity Analysis
The operations in neural networks can be framed in terms of matrix multiplication, allowing us to analyze the complexity of these methods. We report the FLOP following RigL’s computation method in Table 2. If we examine sparse matrix with $pmn$ non-zeros, then theoretically the inference cost of EAST without weight sharing and other standard DST methods is just the cost of the sparse matrix, $\mathcal{O}(pmn)$, and their storage requirement is just $\mathcal{O}(PN)$. With weight sharing, EAST maintains the theoretical network size $\mathcal{O}(PN)$ while physically storing $\mathcal{O}(N-N_{s})$ parameters, resulting in built-in compression on GPU’s without native support for irregular sparsity patterns. It reduces the parameter count including zeros from 23.5M to 13.9M, but slightly increases the theoretical inference FLOP count due to the reuse of learnable parameters. This trade-off potentially offers practical benefits in memory-constrained devices.
While DCTpS achieves higher accuracy than standard DST methods at extreme sparsity levels, it comes with substantial computational overhead that limits its practical deployment benefits. DCTpS uses a dense DCT matrix, which theoretically could have a complexity of $\mathcal{O}(qlogq)$. We found that its theoretical inference FLOPs are about 32X higher than the standard DST at 99.95% and 150X higher at 99.99%. This is because the dense DCT computation dominates at extreme sparsities, becoming an increasingly significant bottleneck as sparsity increases. This additional computational overhead directly explains the inference throughput and latency differences observed in our experiments.
4.2 Ablation Studies
Isolating DyReLU and Weight Sharing. We first augment the standard RigL pipeline with only DyReLU phasing, replacing standard ReLU with DyReLU at initialization and gradually transitioning back to ReLU before the first learning rate schedule update. Weight sharing and cyclic sparsity are excluded in this isolated experiment. The results, summarized in Table 4 demonstrate a clear performance boost when utilizing DyReLU phasing, particularly at higher sparsity levels. At 99.99%, RigL completely loses its expressive capacity. However, as soon as DyReLU Phasing is used as a drop-in replacement for ReLU, it immediately starts converging. We observe that not only is training improved while DyReLU is in effect, but the performance continues to improve even after DyReLU has been completely phased out.
Model Method 99.95% Sparsity 99.99% Sparsity
ResNet-34 RigL w/ ReLU 81.47 10.03
RigL w/ DyReLU 81.80 54.58
RigL w/ WS 81.96 62.94
ResNet-50 RigL w/ ReLU 30.14 10.00
RigL w/ DyReLU 69.57 53.32
RigL w/ WS 82.02 62.71
Table 4: Ablating DyReLU phasing on DST with evaluation at different sparsities. All settings are kept the same except the use of DyReLU phasing vs standard ReLU. Experiments were run on CIFAR-10. Results show improved performance with DyReLU Phasing, especially at higher sparsity levels.
Figure 5: Gradient flow analysis. The top row compares gradient with and without DyReLU. The bottom row compares gradient with and without weight sharing.
Improving gradient flow during early training has been shown to be critical for successfully training sparse networks. Evci et al. [5] demonstrated that methods with better gradient flow, particularly during initialization and early training, consistently achieved superior performance. Building on this insight, we investigated the effect of DyReLU phasing on gradient flow in addition to measuring accuracy.
To analyze this, we computed the sum of the gradient norm of all layers, and compared the gradient flow during training between models using DyReLU phasing against those with standard ReLU. As shown in the top row of Figure 5, models using only standard ReLU have zero gradient norms at 99.99% sparsity and this is reflected in their collapsed accuracy. On the other hand, the networks with DyReLU phasing receive high gradient flow in the beginning where $\beta=1$, during phasing where $10<\beta<$ 1, and maintains healthy gradient flow when $\beta=0$. Results suggest that the network retains some of the gradient-boosting benefits initially provided by DyReLU, We conjecture that this is because DyReLU amplifies important features in the beginning and helps the network settle into a configuration that preserves gradients even after switching to standard ReLU. Gradient flow especially during early stages is a strong indicator of the DST’s success, and helps explain the performance gap between the methods. This supports our theoretical findings in the theorem, confirming that DyReLU phasing indeed provides greater gradient flow than the standard ReLU.
Similarly, we isolate the effect of weight sharing by adding only this mechanism to the RigL pipeline, keeping all other settings the same. Once again, gradient flow has been restored and we see a dramatic increase in gradient flow, as shown in the bottom row of Figure 5. This supports our hypothesis that creating additional gradient paths through weight sharing can help maintain stable training at extreme sparsities.
Effect of Cyclic Sparsity. Cyclic sparsity provides a flexible framework that allows networks to explore a wider parameter space. We perform an ablation study on the effect of cyclic sparsities using ResNet-34 and CIFAR-100 with a target sparsity of 99.95%. As shown in Figure 6, different cyclic schedules - whether starting from high sparsity and decreasing, or from low sparsity and increasing, consistently outperform the static sparsity (red) across all configurations. While each schedule can be tuned differently (e.g., altitude, length of cycle, etc.), we consistently observe improved parameter exploration and generalization. We conclude that having some form of cyclic sparsity may improve the performance over static sparsity. However, the large number of possible configurations makes systematic experimentation challenging. Future work will focus on tuning these schedules to identify optimal schedules for different datasets and architectures.
Figure 6: Ablation study using CIFAR-100 and ResNet-34. Impact of cyclic sparsity on density (left) and validation accuracy (right) compared to static sparsity. All runs are performed using the same settings except cyclic sparsity schedule.
5 Conclusion
In this paper, we introduced EAST, a new DST approach focusing on extreme sparsities that incorporates DyReLU phasing, weight sharing, and cyclic sparsity scheduling. Through extensive empirical evaluation on CIFAR-10, CIFAR-100 and ImageNet using ResNet-34 and ResNet-50, we demonstrated EAST’s ability to achieve competitive performance at extreme sparsities, surpassing existing methods by a large margin. Each component of EAST individually contributes to maintaining robust gradient flow and optimizing parameter exploration at extreme sparsity, while their combined effect can further enhance performance stability. Our results highlight the potential for practical applicability in resource-constrained environments, establishing it as a scalable and effective method for highly compressed networks. This work opens up future research to optimize models at extreme sparsities, an area often overlooked in previous studies.
Acknowledgments. Our work was supported by the UKRI EPSRC funding council (EP/V042270/1), a University of Aberdeen PhD studentship, the HPC facility ”Maxwell”, and the Interdisciplinary Institute.
Supplementary Material
Section A: Preliminaries
A.1 Dynamic Sparse Training
Let $\theta\in\mathbb{R}^{N}$ represent the set of all network weights, and let $\mathcal{M}\in\{0,1\}^{N}$ be a binary mask that defines the active connections in the sparse model:
$\theta_{\mathcal{M}}=\theta\circ\mathcal{M}$
where $\circ$ denotes element-wise multiplication. The sparsity $s$ of the network is the fraction of pruned weights, defined as:
$s=1-\frac{\|\mathcal{M}\|_{0}}{\|\theta\|_{0}}$
where $\|\mathcal{M}\|_{0}$ is the number of active (non-zero) weights in the mask $\mathcal{M}$ and $\|\theta\|_{0}$ is the total number of weights in a dense network. We follow the same criteria for updating connections used in RigL. First, we prune by magnitude, where the smallest magnitude weights are pruned. When pruning, the top-K operation selects the weights with the smallest absolute values:
$\mathcal{M}_{\text{prune}}=\text{ArgTopK}(-|\theta_{\mathcal{M}}|,(1-s)\|\theta\|_{0})$
where $s$ is the sparsity level and $\mathcal{M}_{\text{prune}}$ is the mask for pruned weights. We also adopt the gradient-based growth mechanism from RigL for regrowing weights, where the pruned weights with the largest gradients are reintroduced:
$\mathcal{M}_{\text{grow}}=\text{ArgTopK}(|\nabla_{\theta}L|,s\|\theta\|_{0})$
where $\nabla_{\theta}L$ denotes the gradient of the loss function with respect to the weights. This ensures that important connections, as indicated by their gradients, are restored during training.
A.2 Dynamic ReLU
The DyReLU activation introduces adaptive piecewise linear activation functions that adjust to the input. Given input tensor $x=\{x_{c}\}_{c=1}^{C}$, where $C$ is the number of channels, it is defined as:
$y_{c}=f_{\theta(x)}(x_{c})=\max_{1\leq k\leq K}\left\{a_{k}^{c}(x)x_{c}+b_{k}^{c}(x)\right\}$
where K is the number of functions, and k is the index. $a_{k}^{c}(x)$ and $b_{k}^{c}(x)$ are dynamically generated by a hyper-function $\theta(x)$ based on the input $x$. These dynamic coefficients allow DyReLU to provide more flexibility during training, and in our case enable better parameter exploration.
A.3 Phasing out DyReLU
As training progresses, we phase out all DyReLU and gradually replace them with the standard ReLU. We introduce a hyperparameter that regulates the contribution of DyReLU during training. It makes full contribution when it is 1, and gradually decreases to 0 as training progresses, eventually reaching 0 before the first learning rate decay, at which point DyReLU has been completely replaced by ReLU. We find that this helps the network settle into a more stable configuration, as DyReLU is no longer needed once the network has explored the parameter space adequately. This process removes all DyReLU-introduced parameters.
The decay factor for this transition is denoted as $\beta(t)$, which decays linearly over the course of training:
$\beta(t)=1-\frac{t-t_{\text{start}}}{t_{\text{end}}-t_{\text{start}}}$
where $t_{\text{start}}$ and $t_{\text{end}}$ denote the start and end epochs of DyReLU phasing. We choose $t_{end}$ to be the epoch before the first learning rate schedule update in all of our runs. We initialize with:
$f(x;t=0)=f_{DyReLU}(x)$
During the transition, the activation function progressively shifts toward ReLU:
$f(x;t)=\beta(t)\cdot f_{DyReLU}(x)+(1-\beta(t))\cdot f_{ReLU}(x)$
By the end of phasing $t=T_{\text{end}}$, we have:
$f(x;t=T_{\text{end}})=f_{ReLU}(x)$
This smooth transition from a dynamic to static activation function prevents the network from overfitting to a complex activation pattern while still reaping the benefits of DyReLU in the early stages of training.
B. Implementation Details
We share the hyperparameters used from Section 4 in Table B.1.
Table B.1: Experiment hyperparameters used in this paper. Batch Size (BS), Learning Rate (LR), Epochs, Learning Rate Drop (LR Drop), Optimizer with Momentum(Opt), Weight Decay (WD), Sparse Initialization (Init), Cyclic Sparsity-Related hyperparameters - Number of Cycles (n), Length of Cycle (l), Maximum Parameter Multiplier (m), Update Frequency during Cyclic Sparsity ($\Delta T_{cs}$), and Update Frequency for regular DST ($\Delta T_{dst}$).
Model Method Data BS Epochs LR LR Drop Opt WD Init Cyclic $\Delta T_{dst}$
Epochs n l m $\Delta T_{cs}$
ResNet-34, EAST CIFAR-10/100 128 250 0.1 10x[75,150] SGD(0.9) 1.0e-4 ERK 2 50 3 350 4000
ResNet-50 DCT+RigL+EAST CIFAR-10/100 128 200 0.001 - Adam 5.0e-4 ERK 2 50 3 100 100
ResNet-50 EAST ImageNet 128 100 0.1 10x[25,50] SGD(0.9) 1.0e-4 ERK - - - - 4000